OpenStudy (anonymous):

CALC 3!!! fan, medal and best answer Find the volume below the surface z=(sqrt(16-x^2-y^2)) which lies above the cone Z^2=3x^2+3y^2

OpenStudy (anonymous):

first find the volume of the sphere then subtract the volume of the cone, each of these should have there own double integral.

OpenStudy (anonymous):

what would you do to turn this into spherical coordinates

OpenStudy (anonymous):

rho would be 4 theta and phi would be free, so you'd integrate dphi dtheta. this is because the first equation can be expressed as z^2 +y^2 +x^2=4^2

OpenStudy (anonymous):

wait, so what would the integrals be?

OpenStudy (anonymous):

in which coordinate style?

OpenStudy (anonymous):

dxdy

OpenStudy (anonymous):

oops i mean p2sinphidphidtheta

OpenStudy (anonymous):

\[\int\limits_{-4}^{4}\int\limits_{-4}^{4}\sqrt{16-x ^{2}-y ^{2}}dxdy-\int\limits_{-\sqrt(8)}^{\sqrt(8)}\int\limits_{-\sqrt(8)}^{\sqrt(8)} \sqrt{3x^2+3y^2}dxdy\] the sqrt(8) and -Sqrt(8) can be found by setting sqrt(3x^2 + 3y^2)=sqrt(16-x^2-y^2)

OpenStudy (anonymous):

but wouldnt it be easier with spherical? How would you do this problem with spherical coordinates?

OpenStudy (anonymous):

The first integral would be \[\int\limits_{0}^{2\pi}\int\limits_{0}^{\pi}4\rho^2\sin \phi d \phi d \Theta\] but the second integral is actually easier to leave in rectangular coordinates.

OpenStudy (anonymous):

I just realized my boundaries are off on the first integrals i typed, the first integral should both be from -sqrt8 to sqrt8, the one up there actually takes all of the volume outside the cone not just the volume above it.

OpenStudy (anonymous):

the spherical one is off to the upper coordinate of the phi integral should be .7559

OpenStudy (anonymous):

Oh ok. but when I solved it i got ∫(0,pi/3)∫(0,4)∫(0,2pi) p^2 sinphi dtheta dp dphi

OpenStudy (anonymous):

is that not even close?

OpenStudy (anonymous):

I curious about what you did, I see you've got a triple integral which isn't really useful for three space. Phi never exceeds pi as it is the angle from the z axis, the integral from 0-4 drho would yield the volume of the entire sphere, not just that above the cone and o-pi/3 dtheta is a small wedge of that.

OpenStudy (anonymous):

For this problem, my teacher said to use three space. By the way, the phi doesn't exceed theta since it goes from 0 to pi/3 while theta goes from 0 to 2pi.

OpenStudy (anonymous):

You're right about the phi theta bounds, my mistake. A double integral is an analysis of volume in three space much like a single integral is an analysis of two space. The double integral is the way to go here.

OpenStudy (anonymous):

oh hmm. But when you turn something into spherical doesnt it automatically need three integrals?

OpenStudy (anonymous):

cuz the there would be phi, theta and p?

OpenStudy (anonymous):

I took calc 3 last fall, I'm sorry; you are right your initial integral may also have been correct, again it is my fault.

OpenStudy (anonymous):

No no its ok. I usually don't remember stuff after i take a course. Haha.

OpenStudy (anonymous):

I think all you need from your above integral is a 16(rho)cos(phi), which is 16z as z=rho cos(phi).

OpenStudy (anonymous):

is that what i would be integrating?

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/3}\int\limits_{0}^{2\pi}\int\limits_{0}^{4}\rho^2\sin \Phi 16\rho \cos(\phi) d \rho d\]

OpenStudy (anonymous):

the ending there should read \[d \rho d \Theta d \Phi \]

OpenStudy (anonymous):

Ok, i understand. Finally, solving which i think i can do.